ISFS operations for RELAMPAGO

Overview

ISFS supported the Hydromet research for RELAMPAGO and, as such, was operational for the EHOP, scheduled for 1 June 2018 -- 15 April 2019. Due to some delay in shipping ISFS and RAL equipment, the first week of June was the "burn in" period, during which some sensors were not functioning correctly.

Sites

All sites have instrumentation for basic meteorology, radiation, and soil heating and moisture. In addition, seven of the sites have eddy covariance fluxes with a 3D sonic anemometer and H2O/CO2 IRGA, enabling calculation of the entire surface energy balance. Ten of the sites have optical disdrometers (Ott Parsival2), recording the particle size distribution every minute.

It should be noted that, since all site owners generously volunteered their land, the exact tower placement was not determined completely by scientific requirements, so the flow exposure often was not ideal.  Sites were arranged to optimize fetch for North and Southwest winds.  Users are urged to examine the site photographs when interpreting these data.

(This Table isn't quite up-to-date.  However, this ISFSkmz file is very close.)

Number Location Remarks Type Sonic height (m)

Disdrometer?

Latitude Longitude Altitude (m)
s1 Rio Primero   Basic   Y 31° 19.877'S 63° 38.470'W 253
s2 El Tio   Basic   N 31° 22.856'S 62° 48.640'W 124
s3 San Francisco   Flux 5 Y 31° 25.654'S 62° 7.963'W 116
s4 Esperanza School Basic   Y 31° 23.501'S 60° 54.689'W 25
s5 Canada de Gomez Butterfly Basic   Y 32° 48.341'S 61° 26.097'W 97
s6 Marcos Juarez Test plot Flux 6 N 32° 43.423'S 62° 5.625'W 115
s7 Marcos Juarez Old Alfalfa Flux 6 Y 32° 42.947'S 62° 4.478'W 112
s8 Ordonez Causeway Flux 3 Y 32° 48.216'S 62° 57.615'W 147
s9 Ordonez Lake Flux 7 N 32° 50.325'S 62° 53.111'W 133
s10 Laborde Vet Flux 5 Y 33° 9.335'S 62° 49.405'W 123
s11 La Carlota 4 posts Basic   N 33° 18.288'S 63° 21.312'W 157
s12 Alejandro Roca Dog and son Basic   N 33° 24.199'S 63° 38.206'W 191
s13 Rio Cuarto   Basic   Y 32° 58.048'S 64° 39.136'W 646
s14 Berrotaran Met Observer Flux 3 Y 32° 28.334'S 64° 23.708'W 631
s15 Pilar SMN Basic   Y 31° 40.078'S 63° 52.924'W 335

 

Communications/Data Flow

Two sites, s11 and s13 were found to have poor cell phone coverage during site survey and will use satellite modems for real-time data transmission and remote login. The rest of the sites will use cell phone modems. Through these connections, 5-minute average statistics computed on site will be transmitted to Boulder, where they will be assembled into one data file and used to create data plots in near-real time. Raw sample data will be archived onsite and downloaded manually during site visits every 2 months. If the cell modem data bandwidth permits, raw sample data may be transmitted in real time from the sites using them.

Site servicing

To avoid excessive staff travel time, site visits were scheduled for every two months.  In between, staff from the University of Cordoba generously volunteered their time to fix problems as they occurred.  Nevertheless, this approach resulted in some sensor data outages for several weeks.

All sensor-related events:

s9: The solar panels blew down during a high wind 11/12 Jun.  On 14 Jun, the flux sensors were powered down remotely, but the station still died that night.  On 21 Jun, one panel was reconnected to restart the station.  Both panels were reconnected on 18 Jul.

s4: The radiometer tripod was knocked down and soils dug up by cattle on 14 Jun.  On 26 Jun the station went down completely when all usb devices were stolen from the station.  On 24 Jul the station was mostly restored.

s12: On 14 Jul the mote cable was found disconnected and plugged back in.

s13: On 19 Jul, station found down and restarted.  Mote and 2D also found not working, so changed DSM ports.

s10: On 20 Jul, station restarted after dying after visit on 14 Jul.

s15: On 28 Jul, the TRH died.  Did not respond to usual fixes on 19 Oct and was replaced on 21 Oct.

s5: On 7 Sep, the mote was bad.

s12: On 7 Sep, the mote was not reporting.

s10: On 17 Oct, the TRH fan wasn't working, so the TRH was replaced.

s12: On 18 Oct, the soils and radiometer had been damaged by livestock.  The radiometer and Tsoil were restored.

s2: On 19 Oct, the tipping bucket was resetting, but appeared to fix itself on the same day.  Replaced on 11 Nov.

s5: On 8 Nov, TP01 found out of the ground and was replaced.

s6: On 8 Nov, reconnected the Qsoil.

s7: On 10 Nov, replaced Tsoil.

s4: On 10 Nov, cleared out clogs in tipping bucket.

s14: On 12 Nov, PTB was bad and was replaced.  Qsoil and TP01 were disconnected and were completely replaced.

s9: On 14 Nov, Tsoil was found to be exposed, so reburied.

s12: On 14 Nov, tipping bucket was clogged, so cleaned.  Mote was rebooting, so replaced.  Qsoil, Gsoil, TP01 were removed by livestock, so replaced. Rad stand moved a bit.  On 15 Nov, changed TP01 wisard board to get it going, so calibration off after that.

s11: On 14 Nov, tipping bucket was clogged, so cleaned.  Soil and radiometer cables unplugged, but working once reconnected.

s1: On 16 Nov, mote rebooting.

s2: On 16 Nov, Qsoil not reporting.  Found chewed and replaced on 18 Nov.

s4: On 16 Nov, TP01 reporting bad.

s9: On 16 Nov, Tcase is high -- perhaps bad connection internally.  Logbook comment that day suggests a fix.  NR01 replaced 17 Nov.

s7: On 17 Nov, mote cable reconnected.

s6: On 17 Nov, many mote cables were reconnected and mote replaced due to bad port.

s4: On 22 Nov, NR01 died. Reconnected 13 Jan.

s5: On 29 Nov, mote mostly dead.  Cables reconnected 19 Dec, then cable replaced 20 Jan.

s7: On 29 Nov, TP01 found mostly dead since Oct.  Tsoil bad since 18 Dec.  Qsoil bad since 7 Dec.  All fixed on 15 Jan.

s6: On 14 Dec, Tsoil reading 0.  Replaced 14 Jan.

s9: On 9 Jan, Tsoil and TP01 not reporting.

s13: On 30 Dec, NR01 stopped reporting.  Replaced and tripod releveled 17 Jan.

s9: On 24 Nov, mote mostly bad.  Died completely on 28 Nov.  Fixed by disconnecting NR01 on 19 Dec.  Still not reconnected on 25 Dec. 17 Jan mote replaced, Gsoil, TP01, NR01 back up.

s15: On 27 Nov, mote mostly bad.  Sensors reconnected 18 Dec.

s12: On 27 Nov, suspect that Gsoil is upside down? (have later found some Gsoils wired backwards).

s1: On 14 Dec, TRH dead.  Fixed by remote power cycle on 10 Jan.

s3: On 21 Dec, TP01 dead. Replaced 12 Jan.

s9: On 25 Dec, Qsoil reporting low.

s14: On 10 Jan, Tsoil bad.  Mote rebooting intermittently. Tsoil replaced 17 Jan.

s15: On 11 Jan, NR01 releveled.

s1: 11 Jan, tipping bucket clogged

s3: 12 Jan, tipping clogged.

s5: 14 Jan, tipping clogged.

s7: 15 Jan, tipping clogged.

s8: 15 Jan, tipping clogged; soil sensors exposed; TRH bad

s10: 15 Jan, tipping clogged.

s13: 17 Jan, reburied Tsoil; 

s10: 20 Jan, Tsoil replaced.

s15: 11 Mar, Rad releveled (was off); tipping cleaned

s1: 12 Mar, Rad releveled (was off); TRH housing replaced; tipping cleaned

s2: 12 Mar, tipping cleaned

s14: 13 Mar, tipping cleaned

s9: 14 Mar, tipping clogged

s12: 14 Mar, tipping clogged

s11: 14 Mar, tipping clogged; TRH SHT replaced

s4: 15 Mar, tipping cleaned

s5: 15 Mar, tipping cleaned

s3: 15 Mar, tipping PIC replaced

s8 18 Mar, replaced PTB210 with 220 in different port

s1: 12 Apr, mote not reporting

s5: 12 Apr, TP01 not reporting

s7: 12 Apr, Tsoil all negative

s8: 12 Apr, all USB devices down

s9: 12 Apr, TP01 odd

s11: 12 Apr, all USB sensors down

s14: 12 Apr, Tsoil not reporting

 

Data Quality Processing

Sonic anemometers/H2O/CO2 gas analyzers

Sonic anemometers and H2O/CO2 gas analyzers were deployed at 7 sites, primarily to measure net evapotranspiration.  No problems were noted during operations.

2D sonic anemometers

No problems noted during operations.

Temperature/Relative Humidity sensors

Occasional issues with fans or sensors dying that were replaced.

Barometers

PTB210 barometers were used at all sites to measure the station pressure.  One failed at s14 and was replaced.  Another failed at s8 and was replaced with a PTB220.

Radiometers

Several NR01s died during the project and were replaced.  These were cleaned during service visits, but of course also were rinsed by rain events.  Several of the tripods they were mounted on were noted to have shifted position (not level) during operations.  We'll note these periods here, but keep the data, in case they are still useful (such as to qualitatively determine when there were clouds).

Soil sensors

A set of soil temperature profile sensors, moisture, heat capacity, and heat flux sensors were deployed at each of the 15 sites to measure the soil heat flux term of the surface energy budget.  Several sites had damage due to animal (both wild and domestic) activity as well as exposure due to flooding that washed away soil from the measurement plot.  In addition, several sensors died.  Even more, the Binder cables to these sensors (or the mote that collected their data) became unplugged.  As a result, there are many extended periods (up to 3 months) without soil data.

Temperature: Some probes reported bad values and were replaced. There also may be some calibration issues.  During one set of post-cals, temperatures were off by as much as 0.3 degC.

Moisture: These were initially installed at 2.5cm.  However, GRAINEX data suggested that this is too shallow, allowing an influence (lower moisture) of the surface.  Thus, during the mid-November site visits, Qsoil sensors at all sites were moved to 5cm.  We'll have to evaluate if the calibration changed due to this action.  We need to work up all of the gravimetric soil comparisons.

At least one probe was damaged and was replaced.

Heat capacity: Some probes were replaced.

Heat flux: No issues noted during operations, other than cable, mote, and sensor exposure as noted above.

Precipitation

Tipping buckets were used at all 15 sites.  In addition, Ott optical distrometers were deployed at 10 of these sites, both for comparison and to measure the precipitation size distribution.  Size distribution data are not part of standard ISFS data processing, but are available upon request.

During many station visits, tipping buckets were found to be clogged.  Thus, we expect precipitation amounts from these sensors to be underreported.  In contrast, no isssues were found with the Ott data, so their values are to be preferred, when available.  Test tips often were done as part of station servicing and will be removed during data processing.